OpenStudy (anonymous):

how do you graph: f(x) = 3[[x-2]] like i have no idea what that is.........

OpenStudy (anonymous):

its a quadratic equation it looks like you're graphing

OpenStudy (anonymous):

it is the floor function, aka the greatest integer function

OpenStudy (anonymous):

called a "step function" for obvious reasons.

OpenStudy (anonymous):

so is it like every 3 i make a line

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OH THANKS

OpenStudy (anonymous):

forget about the -2 and the times 3 part for a moment, the greatest integer function says take the greatest integer less than or equal to x so [|5|] = 5 and also [|5.999|]=5

OpenStudy (anonymous):

and \[[\pi]=3\] \[[-\pi]=-4\] etc

OpenStudy (anonymous):

also note that wolfram tells you \[3[x-2]=3[x]-6\]

OpenStudy (anonymous):

okay im confused again...what do you mean?